# MaskCaptioner: Learning to jointly segment and caption object trajectories in videos

### v1

@@ 2 Related work
Open-Vocabulary Video Instance Segmentation (OV-VIS). The OV-VIS task aims to segment, track, and classify objects from an open set of categories in videos (Guo et al., 2025; Wang et al., 2023), using datasets such as LV-VIS (Wang et al., 2023). State-of-the-art methods (Guo et al., 2025; Wang et al., 2023; Fang et al., 2025) commonly use query-based approaches that classify objects by matching visual features with CLIP embeddings (Radford et al., 2021). Methods like OVFormer (Fang et al., 2025) or BriVIS (Cheng et al., 2024) improve this approach by better aligning visual queries with the CLIP space. Unlike these methods focused on CLIP feature matching for classification, our work explores describing objects in natural language (Li et al., 2023a).
@@ 4.2.2 Comparison with the state of the art
We compare MaskCaptioner to state-of-the-art DVOC methods following the standard evaluation protocol on three existing benchmarks : VidSTG in Table 3, VLN in Table 4.2.2 and BenSMOT in Table 4.2.2. DVOC-DS (Zhou et al., 2025) reports results without pretraining, and with their disjoint training training strategy, while OW-VISCaptor (Choudhuri et al., 2024) leverages Mask2former pretrained on COCO for instance segmentation. In Table 3, we include results with the same pretraining data as these methods and show the impact of using our data instead. Pretraining MaskCaptioner on COCO yields better detection and tracking compared to OW-VISCaptor but comparable captioning performance due to the similar captioning head design. Including our data leads to an improvement on all metrics and especially captioning with a 6.7 CapA increase (15% relative improvement). When pretraining on the disjoint DVOC-DS set, we observe a substantial gain in the captioning metric due to the model design. Moreover, we show that using our pretraining sets results in a further performance improvement, while additionally allowing a unified, much faster training (2032 GPU hours (Zhou et al., 2025) vs 208 for our approach). Eventually, our proposed approach can output segmentation masks unlike other methods.
Table 3: Comparison with the state of the art on VidSTG DVOC validation set. All models are finetuned on VidSTG. "temp. agg." refers to including the query temporal aggregation module.
@@
+ temp agg COCO + LVIScap + LV-VIScap 52.7 66.8 71.0 63.0
Table 4: Comparison with state of the art on the VLN DVOC validation set. Mask loss refers to using the segmentation masks in the detection loss. All models are finetuned on VLN.
@@
Table 5: Comparison on the BenSMOT validation set. CapA, and thus CHOTA are not reported on this dataset. All models are finetuned on BenSMOT.
@@
Across all benchmarks, we show that including the temporal aggregation module further improves captioning performance by effectively merging information from multiple video clips, enriching the description of temporally extended actions. We notice that this improvement is only marginal on VLN, which we attribute to the short video lengths in that dataset. Note that we maintain consistent detection models when enabling the temporal aggregation module, keeping the DetA and AssA scores unchanged. Other results employ best-clip captioning.
@@ 4.2.3 Additional ablations
Temporal aggregation. We show the impact of temporal aggregation on DVOC performance in Table 8. Increasing the number of clips for aggregation consistently improves the captioning scores at the expense of higher computational cost. Weighting the clips for aggregation using the detection scores yields a clear improvement over the arithmetic mean. We also observe that performing captioning based on the single clip with best score performs relatively well, which highlights the limits of the complexity of the actions observable in the current benchmarks, e.g. VidSTG.(Zhang et al., 2020).
@@ Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation AD011014323R2 made by GENCI. It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “France 2030" program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI project), the ANR project VideoPredict ANR-21-FAI1-0002- 01, and Paris Île-de-France Région in the frame of the DIM AI4IDF. Cordelia Schmid would like to acknowledge the support by the Körber European Science Prize.
@@ A.4.3 More Implementation details
For all experiments except LV-VIS tuning, we first train the segmentation/detection model, then freeze it and tune the captioning head. Respectively for LVIS/VidSTG/LV-VIS we train for $440k/40k/22k$ for the first stage and $5k/2k/2k$ for the second stage. When tuning pretrained models on VidSTG/VLN/BenSMOT, we train the two stages for $(15k,2k)/(15k,500)/(15k,2k)$ steps, whereas for LV-VIS, we end-to-end tune the model for $2k$ steps. Experiments are run with a batch size of $8$, except when using LVIS+COCO and LV-VIS where we use a batch size of $4$ and for video-level tuning of the captioning head where we use a batch-size of 1. Experiments on LV-VIS are end-to-end trainings with clip-level supervision only. For VidSTG, VLN and BenSMOT experiments we use video-level tuning for captioning with temporal aggregation, with $T_{\textrm{agg}}=8$. For all experiments we train the model with a clips of size $T=2$, and at inference use $T=5/1/1/1$, $T_{\textrm{match}}=1/100/20/40$, $K_{\textrm{match}}=1/7/5/7$ for LV-VIS/VidSTG/VLN/BenSMOT experiments respectively. For the largest dataset (COCO + LVIS) the optimization takes 2 days on 4 H100 GPUs.

### v2

@@ 2 Related work
Open-Vocabulary Video Instance Segmentation (OV-VIS). The OV-VIS task aims to segment, track, and classify objects from an open set of categories in videos (Guo et al., 2025; Wang et al., 2023), using datasets such as LV-VIS (Wang et al., 2023). State-of-the-art methods (Guo et al., 2025; Wang et al., 2023; Fang et al., 2025) commonly use query-based approaches that classify objects by matching visual features with CLIP embeddings (Radford et al., 2021). Methods like OVFormer (Fang et al., 2025) or BriVIS (Cheng et al., 2025) improve this approach by better aligning visual queries with the CLIP space. Unlike these methods focused on CLIP feature matching for classification, our work explores describing objects in natural language (Li et al., 2023a).
@@ 4.2.2 Comparison with the state of the art
We compare MaskCaptioner to state-of-the-art DVOC methods following the standard evaluation protocol on three existing benchmarks : VidSTG in Table 3, VLN in Table 4.2.2 and BenSMOT in Table 4.2.2. DVOC-DS (Zhou et al., 2025) reports results without pretraining, and with their disjoint training training strategy, while OW-VISCaptor (Choudhuri et al., 2024) leverages Mask2former pretrained on COCO for instance segmentation. In Table 3, we include results with the same pretraining data as these methods and show the impact of using our data instead. Pretraining MaskCaptioner on COCO yields better detection and tracking compared to OW-VISCaptor but comparable captioning performance due to the similar captioning head design. Including our data leads to an improvement on all metrics and especially captioning with a 6.7 CapA increase (15% relative improvement). When pretraining on the disjoint DVOC-DS set, we observe a substantial gain in the captioning metric due to the model design. Moreover, we show that using our pretraining sets results in a further performance improvement, while additionally allowing a unified, much faster training (2032 GPU hours (Zhou et al., 2025) vs 208 for our approach). Moreover, our proposed approach can output segmentation masks unlike the other methods.
Table 3: Comparison with the state of the art on VidSTG DVOC validation set. All models are finetuned on VidSTG. "temp. agg." refers to including query temporal aggregation.
@@
+ temp agg COCO + LVIScap + LV-VIScap 55.4 66.8 71.0 64.0
Table 4: Comparison with state of the art on the VLN DVOC validation set. Mask loss refers to using the segmentation masks in the detection loss. All models are finetuned on VLN. "temp. agg." refers to including query temporal aggregation.
@@
Table 5: Comparison on the BenSMOT validation set. CapA, and thus CHOTA are not reported on this dataset. All models are finetuned on BenSMOT. "temp. agg." refers to including query temporal aggregation.
@@
Across all benchmarks, we show that including temporal aggregation further improves captioning performance by effectively merging information from multiple video clips, permitting the description of temporally extended actions. We observe only a marginal improvement on VLN, likely due to the short video lengths of this dataset. Because we do not modify the detection and tracking models when adding temporal aggregation, the DetA and AssA scores are unchanged.
@@ 4.2.3 Additional ablations
Temporal aggregation. We show the impact of temporal aggregation on DVOC performance in Table 8. Increasing the number of clips for aggregation consistently improves the captioning scores at the expense of higher training cost. Due to memory constraints we train with a maximum of 32 clips for aggregation on VidSTG. Weighting the clips for aggregation using the detection scores yields a clear improvement over the arithmetic mean. We also observe that performing captioning based on the single clip with best score performs relatively well, which highlights the limits of the complexity of the actions observable in the current benchmarks, e.g. VidSTG (Zhang et al., 2020).
@@ Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation AD011014323R2 made by GENCI. It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “France 2030" program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI project), the ANR project VideoPredict ANR-21-FAI1-0002- 01. Cordelia Schmid would like to acknowledge the support by the Körber European Science Prize.
@@ A.4.3 More Implementation details
For all experiments except LV-VIS tuning, we first train the segmentation/detection model, then freeze it and tune the captioning head. Respectively for LVIS/VidSTG/LV-VIS we train for $440k/40k/22k$ for the first stage and $5k/2k/2k$ for the second stage. When tuning pretrained models on VidSTG/VLN/BenSMOT, we train the two stages for $(15k,2k)/(15k,500)/(15k,2k)$ steps, whereas for LV-VIS, we end-to-end tune the model for $2k$ steps. Experiments are run with a batch size of $8$, except when using LVIS+COCO and LV-VIS where we use a batch size of $4$ and for video-level tuning of the captioning head where we use a batch-size of 1. Experiments on LV-VIS are end-to-end trainings with clip-level supervision only. For VidSTG/VLN/BenSMOT experiments we use video-level tuning for captioning with temporal aggregation, with $T_{\textrm{agg}}=32/8/8$ respectively. For all experiments we train the model with a clips of size $T=2$, and at inference use $T=5/1/1/1$, $T_{\textrm{match}}=1/100/20/40$, $K_{\textrm{match}}=1/7/5/7$ for LV-VIS/VidSTG/VLN/BenSMOT experiments respectively. For the largest dataset (COCO + LVIS) the optimization takes 2 days on 4 H100 GPUs.